# Compact Language Models via Pruning and Knowledge Distillation

### v1

@@ A.1 Width Pruning
Figure 5: Retraining of searched candidates with 1.8B training tokens.
Figure 6: Comparing depth and width pruning.
Figure 7: Estimated importance of Transformer layers. Left plot shows the perplexity-based importance scores, while the right one BI importance scores.
Loss Tokens MMLU HellaSwag HumanEval
$L_{logits}+L_{is}$ 18.9B 58.0 73.6 26.8
$L_{logits}$ 18.9B 58.3 73.8 26.2
$L_{logits}$ 94B 62.8 79.7 30.4
Table 16: Ablation study for Minitron 8B using the CT blend with and without $L_{is}$, and increased retraining token count.
@@
Table 17: Comparison of loss functions and loss components on Nemotron-3 8B.
@@
Loss components LM loss
$L_{logits}$ 2.155
$L_{logits}$ + $L_{o}$(29:13) 2.145
$L_{logits}$ + $L_{o}$(15:15) + $L_{emb}$ 2.240
$L_{logits}$ + $L_{o}$(23:15) + $L_{emb}$ 2.205
$L_{logits}$ + $L_{o}$(29:15) + $L_{emb}$ 2.203
$L_{logits}$ + $L_{o}$(30:15) + $L_{emb}$ 2.188
$L_{logits}$ + $L_{o}$(31:15) + $L_{emb}$ 2.180
$L_{logits}$ + $L_{o}$(28:12) + $L_{emb}$ 2.141
$L_{logits}$ + $L_{o}$(29:13) + $L_{emb}$ 2.141
$L_{logits}$ + $L_{o}$(29:14) + $L_{emb}$ 2.152
$L_{logits}$ + $L_{o}$(30:14) + $L_{emb}$ 2.150
$L_{logits}$ + $L_{o}$(29:13) + $L_{emb}$ + $L_{i}$(29:13) 2.141
A.2 Retraining with Distillation
@@ A.3 Choice of Losses
A.4 One-shot pruning and distillation vs Iterative pruning and distillation
Figure 8: Accuracy on benchmarks for iterative vs one-shot strategy
One-shot vs iterative within one dimension: In order to understand the best prune-retrain strategy considering a single dimension that can be pruned across the model (depth), we experiment with two different approaches for depth pruning and retraining in order to arrive at the Minitron 8B-Depth-pruned model mentioned above.
@@ A.4 One-shot pruning and distillation vs Iterative pruning and distillation
Figure 9: Comparing depth and width pruning.
A.5 Search
@@
A.6 Compute Resources

### v2

@@ A.1 Width Pruning
Figure 5: LM validation loss curve for retraining of two pruned candidates with (L2, L2) and (L2, Mean) metrics for (batch, sequence) aggregation strategies.
T=1 5.43 1.92
T=2 5.55 1.92
T=4 5.24 1.92
Table 14: Comparison of one-shot importance estimation and pruning vs iterative importance estimation and pruning the embedding dimension from the original size to the target size. LM validation loss is reported before and after lightweight retraining.
A.2 Depth vs. Width Pruning
Figure 6 compares the LM validation curves for width vs. combined depth+width pruning.
Figure 6: Comparison of retraining LM validation loss curves across two pruning choices, width alone vs combined depth and width. We observe a flip in the ranking prior to 200 steps of retraining, showcasing the need for a lightweight retraining phase.
A.3 Retraining with Distillation
@@
Table 15: LM loss comparison for various loss functions and loss components on Nemotron-3 8B. The loss component Llogits alone with forward KLD loss outperforms the rest.
@@ A.4 Choice of Losses
A.5 One-shot vs. Iterative Pruning and Distillation
Loss components LM loss
$L_{logits}$ 2.155
$L_{logits}$ + $L_{o}$(29:13) 2.145
$L_{logits}$ + $L_{o}$(15:15) + $L_{emb}$ 2.240
$L_{logits}$ + $L_{o}$(23:15) + $L_{emb}$ 2.205
$L_{logits}$ + $L_{o}$(29:15) + $L_{emb}$ 2.203
$L_{logits}$ + $L_{o}$(30:15) + $L_{emb}$ 2.188
$L_{logits}$ + $L_{o}$(31:15) + $L_{emb}$ 2.180
$L_{logits}$ + $L_{o}$(28:12) + $L_{emb}$ 2.141
$L_{logits}$ + $L_{o}$(29:13) + $L_{emb}$ 2.141
$L_{logits}$ + $L_{o}$(29:14) + $L_{emb}$ 2.152
$L_{logits}$ + $L_{o}$(30:14) + $L_{emb}$ 2.150
$L_{logits}$ + $L_{o}$(29:13) + $L_{emb}$ + $L_{i}$(29:13) 2.141
Loss Tokens MMLU HellaSwag HumanEval
$L_{logits}+L_{is}$ 18.9B 58.0 73.6 26.8
$L_{logits}$ 18.9B 58.3 73.8 26.2
$L_{logits}$ 94B 62.8 79.7 30.4
Table 18: Ablation study for Minitron 8B with and without the loss component $L_{is}$, and increased retraining token count with $L_{logits}$. Adding $L_{is}$ performs on par with using $L_{logits}$ alone.
One-shot vs Iterative for Importance Estimation and Pruning: refer to Table  14.
One-shot vs Iterative within a Dimension: to understand the best prune-retrain strategy considering a single dimension that can be pruned across the model (depth), we experiment with two different approaches for depth pruning and retraining in order to arrive at the Minitron 8B-Depth-pruned model mentioned above.
@@
Figure 7: Accuracy on MMLU, HellaSwag and HumanEval benchmarks for iterative vs one-shot depth pruning and retraining strategy. One shot pruning and retraining outperforms the iterative approach.
@@
Figure 8: Comparison of LM validation loss curves for one-shot pruning of embeddings, MLP, attention and retraining for 400 steps vs the iterative approach; pruning the embeddings first and retraining for 250 steps, followed by pruning MLP, attention and retraining for additional 150 steps.
A.6 Search
@@
A.7 Single vs. Multi-Phase Training
Table 20 compares the accuracy of single vs. multi-phase training.
Strategy Tokens MMLU HellaSwag PIQA HumanEval
Phase1 + Phase2 113B 54.7 80.3 77.2 25.6
Phase2 only 94B 61.9 80.1 76.7 30.5
A.8 Compute Resources